OpenStudy (anonymous):

How many moles of copper are in 6,000,000 atoms of copper?

OpenStudy (abhisar):

Hello @sandeep140499 ! \(\boxed{\boxed{\boxed{\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}}}}\)

OpenStudy (abhisar):

\[6.022~X~10^{23}\] copper atoms form one mole of it

OpenStudy (abhisar):

So, divide 6,000,000 with it and u will get the no of moles

OpenStudy (kainui):

@sandeep140499 a completely similar but slightly different question is: How many dozen eggs is 36,000 eggs?

OpenStudy (abhisar):

yes very truly said @Kainui

OpenStudy (abhisar):

u just need to apply simple unitary method

OpenStudy (somy):

lol AB i thought you were gonna actually answer xD

OpenStudy (kainui):

I know when I first learned this I didn't really "get" what a mole was. But really it's just another way of grouping things. Hopefully we can help him understand so it doesn't seem so difficult! =)

OpenStudy (somy):

i was talking about @Kainui 's question lmao

OpenStudy (abhisar):

LOL...

OpenStudy (somy):

good job tho :D

OpenStudy (kainui):

I think it's fair to lay it out explicitly with math. \[\Large 6.02 * 10^{23} \ atoms = 1 \ mole\] Now we want to convert from atoms to moles. We are given atoms already, \[\Large 6,000,000 \ atoms\] So to convert, we need to make the unit moles and divide out the atoms, just like we would a number! So we divide from both side and get the number 1. This is the trick. We can multiply anything by 1 and it isn't changed, so that's how conversion factors work. \[\frac{1}{\Large 6.02 * 10^{23}} \ \frac{moles}{atoms} = 1\] Multiply this by the atoms and we get the atoms cancel since anything divided by itself is 1.\[\Large 6,000,000 \ atoms *\frac{1}{\Large 6.02 * 10^{23}} \ \frac{moles}{atoms}=\] \[\Large \frac{6,000,000 }{\Large 6.02 * 10^{23}} \ moles\] I hope that makes it intuitive as to what to do, how to do it, and why.

OpenStudy (somy):

wow awesome job @Kainui !!

OpenStudy (kainui):

Thanks! I don't think the question-asker is even still around but maybe you two got something out of my explanation! =P

OpenStudy (abhisar):

Yes, i learned about eggs

OpenStudy (kainui):

Hahahaha XD

OpenStudy (abhisar):

Thanx All \(\color{green}{\huge\ddot\smile}\)

OpenStudy (kainui):

@Abhisar even better, how many pairs of shoes are in 40,000,000 shoes? =P

OpenStudy (somy):

lol ppl

OpenStudy (abhisar):

aaammm... i think i would have to look wikipedia for this !! :D

OpenStudy (nincompoop):

laughing out loud @Kainui